OpenStudy (anonymous):

Y(1)^3/Y(1)^2=r(i)^3/T Solve for r. Anything in the ( ) is a subscript

OpenStudy (anonymous):

r = Y(1)/ r=(Y(1)/√3y(1)^2)T

OpenStudy (anonymous):

r=(Y(1)/√3y(1)^2)T